OpenStudy (anonymous):

3x + 4y divided by 2

OpenStudy (mysangels):

\[3x+4y \div 2 = 1.5x + 2y\]

OpenStudy (anonymous):

\[\frac{ 3x+4 }{ 2 } = 1.5x+ 2y\]